OpenStudy (anonymous):

When asked to write a point-slope equation for a line passing through the points (-2, -2) and (2, 1), a student wrote y - 1 =-3/4(x-2) . How can she fix this to write the correct equation? A. That is the correct equation, so no changes are necessary. B. Change the minus signs to plus signs. C. Distribute -3/4 to each term in parentheses. D. Change the sign of the slope.

OpenStudy (anonymous):

can u find the slope?

OpenStudy (anonymous):

yea the slope would be -3/4

OpenStudy (anonymous):

u know the equation of that equation (in the question) is as this form: \(y-y_{0} = m(x-x_{0})\) which \( (x_{0} , y_{0})\) is a point (in the question the point (2,1) was used.

OpenStudy (anonymous):

no,use slope formula and find it again...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it would be -2 - 2/-2 -1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

no!

OpenStudy (anonymous):

??

OpenStudy (anonymous):

you have the numbers backwards

OpenStudy (anonymous):

slope formula: \(\Large \frac{ y_{2}-y_{1} }{ x_{2} - x_{1} }\)

OpenStudy (anonymous):